shift summary: day shift 10/14/99

Ashkan_Alavi-Harati (alavi@fnal.gov)
Thu, 14 Oct 1999 15:35:00 -0500

Shift Summary for day shift, 10/14/99

On Shift: Mike Arenton, Ashkan Alavi

Hours of beam available at all intensities: 8.0
Hours of beam available at nominal intensity: 7.5
Hours of E799 data (special runs not included): 7.0
Hours of special runs (laser scans, muon runs...): 0

Accounting of lost time: DAQ problems (our part).

========================================================================
Runs ended this shift:
Run 14835 - BEAM_799 - 19.8M events

Tapes brought to FCC this shift:
(including tapes for online split)

UP*73-75. UPS025
========================================================================
Accesses during shift:

none

========================================================================
Overview of shift:

Good beam for the shift and very smooth 799 data-taking until the damp
server died and it took us an hour to recover. Continued smooth run after
that again.

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
14835 195 07:10 ktevana errors, two KFRENDs.

07:35 onsplit tape UPF018 mounted on ktev2.

220 07:46 ktevana errors, KFREND.
221 07:47 ktevana errors, KFREND.

08:30 SEM has been dropping down to 7.8E12 during
the past hour. Mike calls the control room.

09:12 SEM up to 9.2E12 now.

400 11:46 Beam off, B1 bridge problem. There have been a
number of aborts in the past 0.5 hour.

406 11:54 Beam is back.

12:00 Tape usage and data volume logged on the DAQ
monitoring stopped updating. Called Carmenita, she
is looking at it.

12:13 onsplit tape UPH010 mounted on ktev4.

434 12:35 Tapes are full without warning, the damp server is
probably dead again. We could not stop the run
normally and could not save the online histograms
as expected.

12:45 Change the tapes, drt_fresh_start.

13:10 After drt_fresh_start failure a couple of times at
different stages, Ed tracked down the problem to a
possible bad tape or bad tape drive in ktev4.

13:30 Changed the brand new tape with another brand new
tape and tried drt_fresh_start.

14839 13:40 Sash L. finally managed to start the run after a
couple of KFRENDing.

25 14:11 ktevana errors, KFREND.
26 14:12 ktevana errors, KFREND.
27 14:14 Pipeline error, reset the pipeline.

32 14:20 HCC inefficient, reload HCC.

75 15:20 TeV ramp permit toggle at C3, estimate 5-10 min.

========================================================================
Follow up Notes on Previous Shift

========================================================================
Items to Follow up on Next Shift

Page Jason L. if there is a downtime longer than an hour.